Hi, please help me solve these:
7^x + 7^(x-1) = 8^x
5^x + 0,2^x = 5,2
Thank you

Answer by ankor@dixie-net.com(22740)

7^x + 7^(x-1) = 8^x
has to be x=1, then you have
7 + 1 = 8, remember any no.to the power of 0 = 1
:
5^x + 0.2^x = 5.2
same here x=1
5 + .2 = 5.2